# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

### v1

@@ 3.1 Threat Data: Collection and Preprocessing
Raw Data Collection. Given that threat data is predominantly represented by CVEs [100, 102], we focus on collecting CVE-centered metadata from several vulnerability databases: MITRE-CVE [14], NVD [59], Exploit-DB [62], CWE [16], CAPEC [17], and ATT&CK [18], including (but not limited to) associated threat actors, TTPs, affected systems, attack infrastructures, and strategical mitigation approaches.
@@
To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as GitHub commits, Apache [5], and IBM X-Force [39], incorporating diverse cybersecurity activities and countermeasure strategies. These additional sources enhance the comprehensiveness of our dataset by providing real-world exploit details and remediation insights. Table 8 outlines the data sources and the corresponding metadata collected.
@@
Pre-Training Corpus. Next, we preprocess the collected metadata to construct CVE-centric threat reports, which are used as pre-training corpus. As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality reports and diverse layouts:
@@
    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft [13], Cisco [41], CISA [20], Europol [26], and ENISA [30]. These examples are used as demonstrations $\mathcal{S}$ to guide the generation. For each generation, we select an example report $s\in\mathcal{S}$ and construct a 1-shot prompt, designed to replicate the specific style and layout of $s$.
@@
    • Output Sampling: We adjust two decoding hyperparameters—temperature and top-p—which introduce controlled variability in the outputs. This approach, as studied in [11, 72, 115], helps balance randomness and coherence, ensuring diverse yet meaningful generations.
@@ 1 $D\leftarrow\emptyset$ ;
4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}\oplus s\oplus m$;
@@ 12      $D\leftarrow D\cup\{d\}$;
Table 1: Details of the fine-tuning dataset. Each task targets a specific aspect of CTI analysis, incorporating varying dependencies from prior tasks, and involving different NLP modules: TOM–Topic Modeling; NER–Named Entity Recognition; REL–Relation Extraction; RAG–Retrieval-Augmented Generation (using web resources); REA–(any type of) Reasoning; SUM–Summarization.
@@
❶ Attribution Threat Actor, TTPs, Exploitation Path Initial Evidence
@@
❸ Detection CVE, CWE Identifier Initial Evidence+❶❷
❹ Correlation Related CVE Identifier Initial Evidence+❶❷❸
❺ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
❻ Remediation Patch Tool, Code, Mitigation Advisory Initial Evidence+❶❷❸❹❺
3.2 Curriculum Pre-Training
Next, we pre-train the base LLMs (detailed in Table 9) to embed large-scale cybersecurity knowledge. We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
@@ 3.2 Curriculum Pre-Training
$\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$ (1)
The pacing function in Eq. 1 has three stages: (a) Linear Start – In the early epochs ($t\leq T_{1}$), a progressively increasing portion of $D$ is selected for training. (b) Plateau – In the mid stage ($T_{1}<t\leq T_{2}$), as the LLM $\theta$ acquires foundational cyber threat knowledge, the entire corpus $D$ is typically used for training. (c) Reinforced End – In the final epochs ($t>T_{2}$), the corpus $D$ is augmented with the most recent threat data to reinforce training of emerging cyber threats. This augmentation is balanced by a factor $\beta$. Table 9 shows values of $T$, $T_{1}$, and $T_{2}$.
Pre-Training Objective. Identical to the pre-training objective of Llama [92], we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens. Let $x_{j}$ denote a token in a threat report $d$ with $L_{d}$ tokens, we have $d=\{x_{j}\}_{j=1}^{L_{d}}$. The CLM objective is formulated as:
$\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]$ (2)
Here, the model $\theta$ learns to maximize the likelihood of the next token $x_{j}$ given the preceding context $x_{<j}$, across all tokens in a threat report $d$. This objective enables the model to effectively internalize and represent the knowledge included in the cyber threat corpus $D$.
4 Fine-Tuning, Inference, and Customization
This section details the development of CyLens after pre-training, focusing on task-oriented dataset construction for fine-tuning (section 4.1), modularized inference (section 4.2), and organizational customization (section 4.3).
4.1 Fine-Tuning with Cascading Reasoning
Data Construction. CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis [64, 81]. To equip LLMs with cascading analytical capabilities, we design the fine-tuning dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
@@ 4.1 Fine-Tuning with Cascading Reasoning
Figure 3: An illustration of modularization in CyLens, using the same threat report as described in Example 1.
Table 1 presents the analytical targets for CTI tasks ❶-❻, where each task builds upon the preceding ones. As illustrated in Example 1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
To construct the fine-tuning dataset, we aggregate references from MITRE-CVE, NVD, and ATT&CK. For each reference, we extract available evidence as listed in Table 1 and utilize descriptive reports as the initial evidence set $E_{0}$. We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 1. It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❻. Detailed dataset statistics are provided in Appendix B.
Fine-Tuning. To enable CyLens with cascading reasoning capabilities, we fine-tune it using the dataset $D$ described above. The fine-tuning process follows a standard instruction-tuning objective [106, 44], where the model is trained to generate a sequence of answers conditioned on the preceding context. Let $\theta$ represent the parameters of CyLens. The objective function is defined as:
$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$ (3)
@@
Table 2: Evaluation results for ❶ Attribution tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Threat Actor TTP Exploitation Path
@@
4.2 Modularized Inference
At each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ during inference, we employ various NLP techniques to modularize the analytical process. Table 1 summarizes the NLP techniques used for each task. These techniques are implemented through the following approaches:
Topic Modeling. We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform. Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods [7, 50]. Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❻, and then, based on specific analytical targets, as outlined in Table 1.
Example 2
Topic modeling (Figure 3-(b)).
Using the same example as in 1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$. It then further specifies the analytical target as the threat actor within the threat attribution task.
Named Entity Recognition (NER) and Relation Extraction. Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them. We employ the multi-agent generation approach described in section 3.1, utilizing few-shot prompting, where task-specific entities and their relationships are included as demonstrations within the prompt. This approach facilitates a structured understanding of cybersecurity threats, enhancing the accuracy and contextual relevance of extracted threat intelligence.
Retrieval Augmented Generation (RAG). We implement RAG to retrieve relevant web documents through web screening to enrich threat reports. This is especially useful for zero-
Example 3
NER and relation extraction (Figure 3-(c)).
Using Example 1, the NER identifies key entities "APT28" and "C2 communication" within the threat report. Simultaneously, relation extraction determines the "is associated with" relationship between these entities.
day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated. Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$. To control latency, we implement a cache to store retrieved documents. Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
Reasoning and Summarization. We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required. For example, reasoning is used to analyze severity level (e.g., medium) based on gathered evidence. Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
4.3 Organizational Customization
Although CyLens is developed using massive CTI knowledge, it also enables customization to meet specific organizational interests on particular threat categories, specific products, or assisting an entire organization. Such customization is achieved by further fine-tuning LLMs but specifying the scope of data—for example, concentrating on cyber threats related to a single organization. This targeted customization enhances the relevance of CTI in meeting organizational needs.
We demonstrate CyLens’s potential for tailored customization in section 5.6. Through experiments, we highlight its adaptability and flexibility in addressing diverse CTI challenges across different organizational settings.
5 Evaluation Results
This section systematically evaluates CyLens’s performance on CTI tasks ❶-❻ with the following settings:
Dataset. To ensure practicality, we directly assess threats from NVD [59], MITRE [14], and Exploit-DB [62], categorizing into two groups: (1) Historical – threats reported between 2015 and 2024. (2) Zero-day – newly reported threats from the past two months during our evaluation period in 2025.
For each threat, we compile source references from [14, 59] into structured reports and prompt LLMs for analysis. The prompt structure is illustrated in Figure 4. For different tasks, we modify the system prompt and questions accordingly.
To ensure fairness, no test threats were included in training. Statistics and detailed settings are provided in Appendix C.
@@ 5 Evaluation Results
Baseline. We primarily consider two groups of baselines: (1) State-of-the-art (SOTA) CTI agents, SevenLLM [44] and Lily-Cyber [84], specifically developed for CTI, and (2) Industry-leading general-purpose LLMs, such as ChatGPT and Gemini, which are widely used. Additionally, we investigate the impact of ablation baselines in section 6.1.
@@
Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix C.2.

### v2

@@ 3.2. Threat Corups Construction
Raw Data Collection. Given that threat data is predominantly represented by CVEs (Wikipedia contributors, 2025a, c), we focus on collecting CVE-centered metadata from several threat databases: MITRE-CVE (Corporation, nd), NVD (National Institute of Standards and Technology, 2025), Exploit-DB (Offensive Security, 2025), CWE (Corporation, 2006), CAPEC (Corporation, 2007), and ATT&CK (Corporation, 2013).
@@
To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as Oracle, Red Hat Bugzilla, Adobe Security Bulletin, and the Apache Mailing List, incorporating diverse cybersecurity activities and countermeasure strategies. Table 10 outlines the data sources and the corresponding metadata collected.
@@
Corpus Generation. Next, we curate the collected metadata to construct CVE-centric threat reports, which are then aggregated as threat corpus used for re-training. As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality threat reports and diverse layouts:
@@
    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft (Corporation, 2024), Cisco (Intelligence, 2024), CISA (Cybersecurity and , CISA), Europol (Europol, 2017), and ENISA (for Cybersecurity , ENISA). These examples (or “demos”) $\mathcal{S}$ are used to guide the generation. For each generation, we select an example report $s\in\mathcal{S}$ and construct a 1-shot prompt, designed to replicate the specific style and layout of $s$.
@@
    • Output Sampling: We adjust two decoding hyperparameters, temperature and top-p, which introduce controlled variability in the outputs, thus helps balance randomness and coherence, ensuring diverse yet meaningful generations (Chen et al., 2021; Pearce et al., 2023; Zhu et al., 2024a).
@@ 1 $D\leftarrow\emptyset$ ;
4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}_{\texttt{gen}}\oplus s\oplus m$;
@@ 12      $D\leftarrow D\cup\{d\}$;
Table 1. Base LLMs and re-training settings. “Cp” – Corpus; “Tk” – token. We assign different data volumes each model given their capacities and optimization requirements.
CyLens Base LLM Curriculum Training
# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
@@
Table 2. Details of CTI tasks. Each task targets a specific aspect of CTI analysis, incorporating varying dependencies from prior tasks, and involving different NLP modules: TOM–Topic Modeling; NER–Named Entity Recognition; REL–Relation Extraction; RAG–Retrieval-Augmented Generation (using web resources); REA–(any type of) Reasoning; SUM–Summarization. NLP modules are integrated at inference time to guide CyLens ’s behavior.
@@
❶ Attribution Threat Actor, TTPs, Campaign Initial Evidence
@@
❸ Correlation CVE, CWE Identifier Initial Evidence+❶❷
❹ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
❺ Remediation Patch Tool, Code, Mitigation, Advisory Initial Evidence+❶❷❸❹
3.3. Curriculum Training (Knowledge Infusion)
Next, we re-train base LLMs (detailed in Table 1) to infuse large-scale CTI knowledge. We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
@@ 3.3. Curriculum Training (Knowledge Infusion)
(1) $\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$
The pacing function in Eq. 1 has three stages: (a) Linear Start – In the early epochs ($t\leq T_{1}$), a progressively increasing portion of $D$ is selected for training. (b) Plateau – In the mid stage ($T_{1}<t\leq T_{2}$), as the LLM $\theta$ acquires foundational cyber threat knowledge, the entire corpus $D$ is typically used for training. (c) Reinforced End – In the final epochs ($t>T_{2}$), the corpus $D$ is augmented with the most recent threat data to reinforce training of emerging cyber threats. This augmentation is balanced by a factor $\beta$. Table 1 shows values of $T$, $T_{1}$, $T_{2}$, and $\beta$.
Training Objective. Identical to the re-training objective of Llama (Touvron et al., 2023), we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens. Let $x_{j}$ denote a token in a threat report $d$ with $L_{d}$ tokens, we have $d=\{x_{j}\}_{j=1}^{L_{d}}$. The CLM objective is formulated as:
(2) $\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]+\lambda\|\theta-\theta_{0}\|_{2}^{2}$
Here, the model $\theta$ is trained to maximize the likelihood of each token $x_{j}$ given its preceding context $x_{<j}$, across all tokens in a threat report $d$. To prevent overfitting to the training corpus, we introduce a regularization term $\|\theta-\theta_{0}\|_{2}^{2}$, where $\theta_{0}$ denotes the original pre-trained parameters. This regularization encourages the model to retain its inherent language understanding and reasoning capabilities. The resulting objective function strikes a balance between effective CTI knowledge infusion and preserving the general-purpose strengths of the base LLM.
3.4. Instruction Tuning via Cascading Reasoning
Data Construction. CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis (Oosthoek and Doerr, 2021; Saeed et al., 2023). To equip LLMs with cascading analytical capabilities, we design a dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
@@ 3.4. Instruction Tuning via Cascading Reasoning
$\mathbf{E_{0}}$: A security monitoring system detected suspicious outbound traffic from several workstations within the organization’s network. Traffic analysis indicates command-and-control (C2) communication with IP addresses known to be associated with APT28, with observed indicators of $\cdots$
Step 1: Threat Actor Attribution
$\mathbf{Q_{1}}$: From above, what threat actor is likely responsible?
@@
Step 2: Vulnerability Identification
$\mathbf{Q_{2}}$: What known vulnerabilities (CVEs) are commonly associated with APT28 in recent campaigns?
$\mathbf{A_{2}}$: APT28 has exploited CVE-2023-12345 (Windows Kernel Privilege Escalation) and CVE-2023-67890 (Remote Code Execution via malicious DLL sideloading).
Step 3: Prioritization of Threat Response $\cdots$
Figure 3. An illustration of modularization in CyLens, using the same threat report as described in Example 3.1.
[…]
Modularized inference.
Table 2 presents the analytical targets for CTI tasks ❶-❺, where each task builds upon the preceding ones. As illustrated in Example 3.1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
To construct the dataset, we aggregate references from threat databases (e.g., MITRE and NVD). For each reference, we extract available evidence as listed in Table 2 and utilize descriptive reports as the initial evidence set $E_{0}$. We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 3.1. It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❺. Dataset statistics are provided in Appendix B.
Instruction Tuning. To enable CyLens with cascading reasoning capabilities, we fine-tune it using the dataset $D$ described above. This process follows a standard instruction-tuning objective (Xu et al., 2023; Ji et al., 2024), where the model is trained to generate a sequence of answers conditioned on the preceding context. Let $\theta$ represent the parameters of CyLens. The objective function is defined as:
(3) $\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$
@@
3.5. Modularized Inference
CyLens’s inference also follows the cascading reasoning layout. Each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ employs various NLP functions to modularize the analytical process. Table 2 summarizes the NLP techniques applied to each CTI task. These techniques are implemented through the following approaches:
Topic Modeling. We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform. Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods (Blei et al., 2003; Lee and Seung, 2000). Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❺, and then, based on specific analytical targets, as outlined in Table 2.
Example 0.
Topic modeling (Figure 3-(b)).
Using the same example as in 3.1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$. It then further specifies the analytical target as the threat actor within the threat attribution task.
Named Entity Recognition (NER) and Relation Extraction. Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them. We employ the multi-agent generation as described in section 3.2, using few-shot prompting to instruct entity and relation extraction, where domain-specific entities (e.g., APT group) and their relationships are included as demonstrations within the prompt. We defer detailed implementation in section C.2 and C.3.
Example 0.
NER and relation extraction (Figure 3-(c)).
Using Example 3.1, the NER identifies key entities ”APT28” and ”C2 communication” within the threat report. Simultaneously, relation extraction determines the ”is associated with” relationship between these entities.
Retrieval Augmented Generation (RAG). We implement RAG to retrieve relevant web documents through web screening to enrich threat reports. This is especially useful for zero-day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated. Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$. To control latency, we implement a cache to store retrieved documents. Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
Reasoning and Summarization. We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required. For example, reasoning is used to analyze severity level based on gathered evidence. Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
Additional implementation details (e.g., prompts) of the six NLP modules are provided in appendix C.
4. Experimental Setting
Table 3. Evaluation results for ❶ Attribution tasks, where ”Hist.”–historical threats, ”0-Day”–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Threat Actor TTP Campaign
@@ 4. Experimental Setting
This section outlines our experimental setup for CTI tasks ❶–❺, designed to evaluate CyLens’s scalability and adaptability. The corresponding results are presented in section 5 and 6, respectively.
SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI). Your role is to analyze threat reports, ¡Specify Analytical Target¿, and produce structured threat intelligence analysis grounded in established frameworks (e.g., MITRE ATT&CK). You will follow a structured analytical pipeline with the following steps: Resolve Task Dependencies: Determine the prerequisite CTI tasks and analytical targets based on the input query. Launch NLP Modules: ¡Explain Six NLP Modules¿. USER ¡Report Title¿ ¡Report Excerpt¿ ¡User’s Question¿ ASSISTANT The user’s intent pertains to ¡Topic Modeling¿, which aligns with the prerequisite identification of ¡CTI Tasks¿ and corresponding ¡Analytical Targets¿. Proceeding with a structured inference pipeline, we will now extract and analyze the following components: ¡Key Entities¿, ¡Entity Relations¿, ¡Retrieved Evidence¿, ¡Sources¿, and ¡Synthetic Reasoning¿. Final Response: ¡Summarization¿
Figure 4. Inference Prompt structure. For baselines and CyLens, we prompt them to perform cascading reasoning with the NLP modules defined in Table 2. For tasks with prior dependencies, we instruct LLMs to first complete the prerequisite tasks before proceeding.
Scenarios. To evaluate the scalability and adaptability of CyLens, we consider two experimental scenarios:
    • Broad CTI — Demonstrates CyLens’s scalability in analyzing a wide range of cyber threats across CTI tasks ❶–❺, providing a comprehensive view of the threat landscape.
    • Specialized CTI — Demonstrates CyLens’s adaptability to specific cyber threat groups, such as those targeting particular products or vendors tailored to business needs, or threats within a specific category to support focused interests (e.g., (Mishra et al., 2022; Balasubramanian et al., 2025)).
In both scenarios, CyLens undergoes the same re-training using the threat corpus described in Table 1. However, instruction tuning differs based on the evaluation focus. For the broad CTI setting, we use a diverse instruction tuning dataset covering a wide spectrum of threats, as reported in Table 11. In the Specialized CTI setting, we extract targeted subsets of threats from the broader dataset and independently fine-tune CyLens to emphasize its functions on specific threat groups (e.g., those associated with a particular product, vendor, or threat category).
Dataset. To ensure practical evaluation, we directly collect threats reported from real-world platforms, including Oracle Security Alerts (Oracle Corporation, 2025), Red Hat Bugzilla (Red Hat, Inc., 2025), Adobe Security Bulletin (Adobe Systems Incorporated, 2025), NVD (National Institute of Standards and Technology, 2025), VulDB (VulDB, 2025), Patchstack (Patchstack, 2025), and public mailing lists from Apache (Apache Software Foundation, 2025a), Debian (Debian Project, 2025), openSUSE (openSUSE Project, 2025), and Fedora (Fedora Project, 2025). To avoid data leakage, all threats correlated with these evaluation reports are excluded from both the re-training and instruction tuning phases. We categorize the evaluation threats into two groups: (1) Historical — Threats disclosed between 2015 and 2024. (2) Zero-Day — Newly disclosed threats from the two months immediately preceding our evaluation period in 2025. Table 4 summarizes the statistics of the evaluation dataset.
Table 4. Statistics of evaluation threats (not training) for broad and specialized CTI. ”Mc.”–MacOS, ”Cr.”–Chrome, ”Lk.”–Linux Kernel, ”Ad.”–Adobe, ”Dl.”–Dell, ”Of.”–Overflow
# Threat Broad Product Vendor Category
Mc. Cr. Lk. IBM Ad. Dl. Of. DoS XSS
Historical 26.5K 2.5K 1.7K 4.6K 3.3K 1.8K 0.8K 7.5K 9.6K 12.8K
Zero-Day 3.6K 141 48 176 81 72 33 407 721 2.6K
@@
Baseline. We primarily consider two groups of baselines: (1) SOTA CTI agents, SevenLLM (Ji et al., 2024) and Lily-Cyber (Segolily Labs, 2025) and (2) Industry-leading general-purpose LLMs: ChatGPT-4o/o1, Gemini-Pro, Llama-405B, DeepSeek-R1, and Claude-Opus. Additionally, we investigate the impact of ablation baselines in section 7.1.
@@
Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix D.2.
5. Evaluation Results: Broad CTI